what would be $/₺ rate after a year if expected inflation in Turkey is projected as %18.5 whereas inflation for USA is projected to be %3.5? Today's spot rate

1$= 5.35 ₺
Answer #1

Purchasing Power Parity Theory (PPT) is given by

1+Ih / 1+If = F1/S0

where

Rh - Inflation rate in home country = 18.50%

Rf - Inflation rate in foreign country = 3.50%

F1 - 1 year forward rate =?

S0 - Spot Rate =5.35

Here exchange rate is given in the format of t/$.  Hence, Turkey (first currency) is the home country and USA (second currency) is the foreign country.

1.185/1.035 = F1/5.35

F1/5.35 = 1.14492753623

F1 = 1.14492753623*5.35

= 6.13t/$

= 1/6.13 $/t

= 0.1631 $/t

= 0.16 $/t

PPT states that high Inflation rate in a country is offset by depreciation in the currency of that country. So here t should depreciate.

Here forward rate>spot rate. That is product in the exchange rate, $ is appreciating and t is depreciating, which is in line with PPT.